# shapes.geometric, shapes.multipart, arrows, arrows.meta, positioning, calc, chains, backgrounds Towards Explainable Evaluation Metrics for Machine Translation

### v1

@@ 1 Introduction
The field of evaluation metrics for Natural Language Generation (NLG), especially machine translation (MT) is in a crisis (Marie et al. 2021).111We published an earlier version of this paper (Leiter et al. 2022a) under a different title. Both versions consider the conceptualization of explainable metrics and are overall similar. However, while the old version contains several novel experiments, this new version puts a stronger emphasis on the survey of approaches for the explanation of MT metrics including the latest LLM based approaches. For example, this comprises techniques that return fine-grained error labels or natural language explanations. Despite the development of multiple high-quality evaluation metrics in recent years (Zhao et al. 2019; Zhang et al. 2020a; Rei et al. 2020; Sellam et al. 2020; Yuan et al. 2021), the Natural Language Processing (NLP) community appears hesitant to adopt them for assessing NLG systems (Marie et al. 2021; Gehrmann et al. 2023). Empirical investigations of Marie et al. (2021) indicate that the majority of MT papers relies on surface-level evaluation metrics such as BLEU and METEOR (Papineni et al. 2002; Banerjee and Lavie 2005), which were created two decades ago, a trend that may allegedly have worsened in recent times. These surface-level metrics cannot (even) measure semantic similarity of their inputs and are thus fundamentally flawed, particularly when it comes to assessing the quality of recent state-of-the-art MT systems (e.g. Peyrard 2019; Freitag et al. 2022), raising concerns about the credibility of the scientific field. We argue that the potential reasons for this neglect of recent high-quality metrics include: (i) non-enforcement by reviewers; (ii) easier comparison to previous research, for example, by copying BLEU-based results from tables of related work (potentially a pitfall in itself); (iii) computational inefficiency to run expensive new metrics at large scale; (iv) lack of trust in and transparency of high-quality black box metrics. In this work, we concern ourselves with the last-named reason, and address the explainability of such metrics.
In recent years, explainability has become a crucial research area in AI due to its potential benefits for users, designers, and developers of AI systems Samek et al. (2018); Vaughan and Wallach (2021).222 As of now, there is no universally accepted definition of explainability in the AI community. In this work, we adapt the definition by Barredo Arrieta et al. (2020), which we discuss in §2.2.
@@
For users of the AI systems, explanations help them make more informed decisions (especially in high-stake domains) Sachan et al. (2020); Lertvittayakumjorn et al. (2021), better understand and hence gain trust of the AI systems Pu and Chen (2006); Toreini et al. (2020), and even learn from the AI systems to accomplish tasks more successfully Mac Aodha et al. (2018); Lai et al. (2020). For AI system designers and developers, explanations allow them to identify the problems and weaknesses of the systems Krause et al. (2016); Han et al. (2020), calibrate the confidence of the systems Zhang et al. (2020b), and improve them accordingly Kulesza et al. (2015); Lertvittayakumjorn and Toni (2021).
@@
Explainability is particularly desirable for evaluation metrics. Sai et al. (2022) suggest that explainable NLG metrics should focus on providing more information than just a single score (such as fluency or adequacy). Celikyilmaz et al. (2020) stress the need for explainable evaluation metrics to spot system quality issues and to achieve a higher trust in the evaluation of NLG systems. Explanations indeed play a vital role in building trust for new evaluation metrics.333As an illustrating example, Moosavi et al. (2021) distrust using the metric BERTScore Zhang et al. (2020a) applied to a novel text generation task, as it assigns a score of 0.74 to a nonsensical output, which could be taken as an unreasonably high value. While BERTScore may indeed be unsuitable for their novel task, the score of 0.74 is meaningless here, as evaluation metrics may have arbitrary ranges. In fact, BERTScore typically has a particularly narrow range, so a score of 0.74 even for bad outputs may not be surprising (BERTScore provides a scaling method that solves these issues to some degree). Explainability techniques would be very helpful in preventing such misunderstandings, e.g. by showing that the features BERTScore attends to align with human judgement. For instance, one potential reason that a metric will likely be better accepted by the research community is if the explanations of the scores align well with human reasoning and faithfully reflect its internal decision process (see §2.2).
@@
By contrast, if faithfulness is given and the explanations are counter-intuitive, users and developers might lower their trust and be alerted to take additional actions, such as trying to improve the metrics using insights from the explanations or looking for alternative metrics that are more trustworthy. Furthermore, explainable metrics can be used for other purposes: for example, when a metric produces a low score for a given input, highlighted words (a widely used method for explanation, see §4) in the input are natural candidates for manual post-editing.
@@
This concept paper aims at providing a systematic overview of the existing efforts in explainable MT evaluation metrics and an outlook for promising future research directions. The main focus of this work lies on explainable evaluation metrics for MT, many of our observations and discussions can likely be adapted to other NLG tasks, however.
@@
§5 Future Work: An analysis of underexplored research directions , and a collection of exemplary methods from other NLP domains to illustrate potential future paths. Here, we also discuss the usage of large language models (LLMs), like ChatGPT OpenAI (2023b) and GPT4 OpenAI (2023a).
With this work, we aim to solidify the field of explainable MT metrics and provide guidance for researchers, in specific metric developers, who aim to better understand and explain MT metrics, as well as users that want to employ explainability techniques to explain metric outputs. In the long term, we envision that our work will aid the development of improved MT metrics and thereby help to improve the quality of machine translations. Potentially, explanations can also aid other use cases such as translation selection and semi-automatic labeling (see §2.3). We also hope that our work will more generally inspire explainable NLG metric design.
@@ 2.1 Machine Translation Evaluation Metrics
MT metrics grade machine translated content, the hypothesis, based on ground truth data. We differentiate these metrics along several dimensions. A summary is shown in Table 1. We note that metrics can be categorized along many dimensions (e.g. Sai et al. 2022; Celikyilmaz et al. 2020). Here, we focus on a minimal specification that we will leverage later. Some parts of this section refer to explainability (see definitions in §2.2).
@@
Whether source, reference translation or both are used as ground truth for comparison |
| Granularity |
@@
At which level a metric operates: Word-level, sentence-level, document-level |
| Quality aspect |
What a metric measures: Adequacy, fluency, etc. |
| Learning objective |
How a metric is induced: Regression, ranking, etc. |
@@
Input type We call a metric reference-based if it requires one or multiple human reference translations to compare with the hypothesis as ground truth, which can be seen as a form of supervision (e.g. Papineni et al. 2002; Zhao et al. 2019; Rei et al. 2022a). Reference-free metrics do not require a reference translation to grade the hypothesis (e.g. Zhao et al. 2020; Ranasinghe et al. 2020b; Belouadi and Eger 2023). Instead, they directly compare the source to the hypothesis. In the literature, reference-free MT evaluation is sometimes also referred to as “reference-less” (e.g. Mathur et al. 2020) or “quality estimation” (e.g. Zerva et al. 2022). This dimension is important for explainable MT evaluation, as explainability techniques might need to consider which resources are available.
@@
Granularity Translation quality can be evaluated at different levels of granularity: word-level, sentence-level and document-level. The majority of metrics for MT return a single sentence-level score for each input sentence (e.g. Zhao et al. 2019; Zhang et al. 2020a). Beyond individual sentences, a metric may also score whole documents (multiple sentences) (Jiang et al. 2022; Zhao et al. 2023). In the MT community, metrics that evaluate translations at the word-level (for example, whether individual words are correct or not) are also common (Turchi et al. 2014; Shenoy et al. 2021). Recently, there has been a tendency to compare metrics to human scores derived from fine-grained error annotations — multidimensional quality metrics (MQM) (Lommel et al. 2014) — as they tend to correspond better to human professional translators (Freitag et al. 2021a). Metrics of higher granularity (e.g. word-level) provide more explanatory value than those of lower granularity, as they provide more details on which parts of the input might be translated incorrectly (e.g. Leiter 2021; Fomicheva et al. 2021, 2022). Most techniques we describe in §4 target the explainability of sentence-level metrics.
@@
Quality aspect This refers to the properties a metric evaluates and measures, for example, human assigned adequacy (via direct assessment (Graham et al. 2016) obtained from crowd-workers), fluency, or other aspects (such as relevance, informativeness, or correctness, mostly in other NLG fields) (e.g. Sai et al. 2022; Celikyilmaz et al. 2020). We use this dimension in §4, as metrics that report multiple quality aspects provide more information on specific translations, i.e., they have more explanatory value for potential users.
@@
Learning objective Yuan et al. (2021) describe a further distinction based on the task a metric is designed to solve (or how a metric is implemented) to achieve high correlation with human judgments. They identify the three tasks below.444Chen and Eger (2023) argue for metrics induced from natural language inference (NLI) as a conceptual framework.
@@
Supervised Regression/Ranking: These models use supervised regression to train a metric to predict continuous values that directly correlate with human judgments from hypothesis and ground truth tokens. Alternatively, a metric may be trained to rank hypothesis sentences. For example, Rei et al. (2020) propose a ranking setup in which they minimize a loss that becomes smaller the greater the distance between embedding representations of the two sentences become. Especially in the MT community, metrics trained on human scores (e.g. COMET; Rei et al. 2022a) have become more and more dominant, while Belouadi and Eger (2023) argue that unsupervised metrics are more widely applicable.
@@
Text Generation: Some metrics use the probability that a sentence (e.g., a translation) is created by a text generation system, given an input sentence (e.g., a reference sentence), as sentence-level score. Yuan et al. (2021) place the metric PRISM (Thompson and Post 2020) and their own metric BARTScore in this setup.
@@
MT metrics also have multiple use cases; see Figure 1 (in the following, preceding numbers correspond to this figure). (1) One use case is to rate a single translated text, for example, in a scenario where a human user wants to check their own human555Note, however, that humans might make different types of translation errors (Specia et al. 2018). Therefore, metric performance first needs to be evaluated for this use case. translation or a translation they received by an MT model before using it in a downstream task (e.g. Murgolo et al. 2022; Zouhar et al. 2023). (2) The standard use case is to rate a large number of translations to compare MT systems (e.g. Freitag et al. 2022). (3) Another use case is to employ metrics in the training process of MT models, e.g. using reinforcement learning (e.g. Wu et al. 2018) or optimizing by computing derivatives with Gumbel-softmax (e.g. Jauregi Unanue et al. 2021). (4) The fourth use case is to use the metrics in the decoding process of machine translation systems for reranking or with minimum Bayes risk decoding Fernandes et al. (2022a).
The use cases can have different audiences and explainability might play a different role to them (see 2.2). For example, anyone might be a user that wants to rate a translation, but usually only domain experts want to rate a translation model.
@@ 2.2 Explainability
Explainability & interpretability Prior work has noted that the terms interpretability and explainability vary in definition and are sometimes used ambiguously throughout the literature (Barredo Arrieta et al. 2020; Jacovi and Goldberg 2020; Vilone and Longo 2021). In this paper, we follow the notion of passive interpretability and active explainability (for example, Barredo Arrieta et al. 2020). In particular, we adapt the definition of Barredo Arrieta et al. (2020) as follows:
@@
This definition varies from Barredo Arrieta et al. (2020) in two aspects: (1) It includes external explainers (also called post-hoc methods) and (2) it includes explanations of the metric’s output. We note that (2) is different from explaining a metric, as output explanations may not provide insights into how a model works. Still, we include (2) in our explainability definition for two reasons. One is that there have been two shared tasks and other existing works on metric explainability that consider the plausibility aspect of explanations only (see the following paragraph on faithfulness and plausibility) (e.g. Fomicheva et al. 2020, 2021; Zerva et al. 2022; Xu et al. 2023). Thereby, explanations these works consider might be explanations of the output that do not actually describe model internals. We want to accommodate these works that do not check faithfulness in our definition. The other reason is that we see use cases of output explanations of strong metrics, i.e., plausible-only explanations (see the following paragraph on faithfulness and plausibility).
@@
In contrast, a metric is interpretable if it can be passively understood by humans (for example, decision trees and k-nearest neighbors) (Barredo Arrieta et al. 2020). Modern MT metrics are mostly based on embeddings and black box language models, making them non-interpretable. The papers we categorize in this work generate explanations for MT metrics and their outputs, i.e. they consider explainability.
[set/.style = text opacity = 1]
[…]
Faithfulness & plausibility An explanation is faithful if it accurately represents the reasoning process underlying a model’s prediction(s) (Jacovi and Goldberg 2020). In contrast, an explanation is plausible if it convinces its audience that a model’s prediction is correct (Jacovi and Goldberg 2021), i.e., if it aligns with human judgment. Both properties are not necessarily dependent on each other (Jacovi and Goldberg 2020). For example, an explanation of an MT metric score could make sense to a human without reflecting the metrics internal details and vice-versa. A degree of faithfulness is indispensable to all explainability works that actually want to reason about a model’s decision processes. Plausibility, on the other hand, might convince explainees of decision processes that do not reflect the model’s internals correctly, or incorrect model predictions. As such, optimizing for plausible explanations might be unethical and lead to deceiving outcomes (Jacovi and Goldberg 2021; Jin et al. 2023). However, in our taxonomy, we find that most current works on explainable MT metrics do not consider faithfulness and only evaluate plausibility (see §4), often leaving unknown whether and to which degree the model’s internal process is accurately described by explanations. While this is a decisive current shortcoming, we argue that explanations that may be only plausible can still benefit certain audiences and use cases of MT metrics that do not target understanding the metric: (1) We assume that current high-quality metrics already work well to judge translation quality in many settings (Freitag et al. 2022; Zerva et al. 2022). Explainers that are proven to have a high plausibility to human experts, as for example measured by Fomicheva et al. (2021); Zerva et al. (2022), will likely provide strong reasoning about the translation quality when provided with a correct score by such metrics. We caution that this process will have two points of failure: metric and explainer, hence the overall rate of failures will be higher than for each of them individually. Still, we predict that this process will work reasonably well to foster the goals of metric accessibility (see §2.3), as long as users are aware of possible failures (see future work §5), similar to the problem of hallucinations in LLMs (Ji et al. 2023b). (2) Language experts should be able to discern incorrect explanations. For them, explanations of quality scores (produced by high quality metrics) that strive to be plausible could provide means to faster evaluating and (if necessary) correcting translations, as encompassed by the goals of metric accessibility and semi-automatic labeling (see §2.3). (3) LLM’s have been shown to provide better results when being prompted to explain them (Wei et al. 2022). Here, more plausible explanations might lead to better scores and vice-versa (see §5). (4) Similar to (3), language models sometimes benefit from multitask learning (Radford et al. 2019), so employing automatically generated explanations might assist training better metrics. Note that (1) and (2) correspond to the first metric use case in Figure 1. To summarize, plausible but not faithful explanations could still be useful either to human users or intrinsically to a metric itself in several scenarios.
Fairness & robustness In Figure 2, we show the relationship of our definition of explainability to robustness and fairness. Robustness measures how well a model (here, a metric) can address execution errors, erroneous and malicious inputs and unseen data (e.g Li et al. 2023). In contrast, Li et al. (2023) define fairness as the goal that a model should not have biases that result in unfair treatment for certain groups of people. As umbrella term for these concepts, we use responsible AI Barredo Arrieta et al. (2020).666There are many related and partly conflicting definitions and terms in explainable AI research (e.g. Vilone and Longo 2021) and the given definitions are not yet firmly established in the literature. For example, Sharma et al. (2020) define explainability, robustness and fairness as exclusive properties. The main focus of our work is to build a taxonomy of methods for explainable MT metrics and the chosen definitions allow us to reasonably relate existing work. Barredo Arrieta et al. (2020) define responsible AI as a concept that encourages the systematic adoption of several AI principles, while also setting a focus on accountability and privacy. Methods that test for robustness and fairness can all be grouped into the given explainability definition, as their fulfillment requires explanations that will also help certain audiences to understand a model’s behavior better Barredo Arrieta et al. (2020).777Barredo Arrieta et al. (2020) group robustness under the term confidence. In the MT metrics literature, this overlap in terminology is apparent for techniques that check how robust metrics are to input perturbations. Here, some works frame these as explainability techniques (e.g. Sai et al. 2021; Karpinska et al. 2022), while others frame them as robustness tests (e.g. He et al. 2022; Chen and Eger 2023); see §4. In Figure 2, we only let robustness and fairness half overlap with explainability to highlight that the explanatory value increases with fine-grained explanations. E.g., robustness could be measured coarse-grained with a single score, such as the minimal number of adversarial perturbations required to successfully attack a metric, or fine-grained, by using predefined perturbation templates (e.g. Sai et al. 2021) or attackers that will attack with perturbations of specific properties (e.g. Chen and Eger 2023). While a single score can also explain whether metric behavior is robust, susceptibility scores to various properties allow to assess metrics in greater detail. The same is true for fairness (Sun et al. 2022). Another reason that we do not fully include robustness and fairness in the explainability definition is that explainability results from testing them, but vice-versa robustness and fairness are usually not given just by having an explainable model.
We note that common explanation types in the evaluation of robustness and fairness are counterfactual explanations, i.e., minimal perturbed input samples that change the output of a model to demonstrate its behavior Wachter et al. (2017); Freiesleben (2021). For fairness evaluation, counterfactuals from domains that carry a risk of being biased are used Sharma et al. (2020); Sun et al. (2022). For robustness, existing works check whether metric scores change as expected, when presented with counterfactuals in contrast to the original hypotheses (e.g. Sai et al. 2021; Chen and Eger 2023); see §4 for more details. Some works frame these counterfactuals as adversarial attacks (e.g. Sharma et al. 2020; Chen and Eger 2023), others as diagnostic tests(ets) (e.g. Karpinska et al. 2022) or challenge sets (e.g. Freitag et al. 2022).
@@
understanding and model-understanding (e.g. Gehrmann et al. 2020; Holzinger et al. 2022).
@@
explaining general properties of a model that steer its decisions. Often, decision-understanding is also referred to as local and model-understanding as global (e.g. Lipton 2016; Guidotti et al. 2018; Doshi-Velez and Kim 2017; Danilevsky et al. 2020).

### v2

# Towards Explainable Evaluation Metrics for Machine Translation

@@ Abstract
Keywords: evaluation metrics, explainability, interpretability, machine translation, machine translation evaluation
@@ 1 Introduction
The field of evaluation metrics for Natural Language Generation (NLG), especially machine translation (MT) is in a crisis (Marie et al., 2021). Despite the development of multiple high-quality evaluation metrics in recent years (e.g., Zhao et al., 2019; Zhang et al., 2020a; Rei et al., 2020; Sellam et al., 2020; Yuan et al., 2021; Rei et al., 2023a; Kocmi and Federmann, 2023a), the Natural Language Processing (NLP) community appears hesitant to adopt them for assessing NLG systems (Marie et al., 2021; Gehrmann et al., 2023). Empirical investigations of Marie et al. (2021) indicate that the majority of MT papers relies on surface-level evaluation metrics such as BLEU and METEOR (Papineni et al., 2002; Banerjee and Lavie, 2005), which were created two decades ago, a trend that may allegedly have worsened in recent times. These surface-level metrics cannot (even) measure semantic similarity of their inputs and are thus fundamentally flawed, particularly when it comes to assessing the quality of recent state-of-the-art MT systems (e.g., Peyrard, 2019; Freitag et al., 2022), raising concerns about the credibility of the scientific field. We argue that the potential reasons for this neglect of recent high-quality metrics include: (i) non-enforcement by reviewers; (ii) easier comparison to previous research, for example, by copying BLEU-based results from tables of related work (potentially a pitfall in itself); (iii) computational inefficiency to run expensive new metrics at large scale; (iv) lack of trust in and transparency of high-quality black box metrics. In this work, we concern ourselves with the last-named reason, and address the explainability of such metrics.
@@
In recent years, explainability has become a crucial research area in AI due to its potential benefits for users, designers, and developers of AI systems Samek et al. (2018); Vaughan and Wallach (2021).111As of now, there is no universally accepted definition of explainability in the AI community. In this work, we adapt the definition by Barredo Arrieta et al. (2020), which we discuss in §2.2. For users of the AI systems, explanations help them make more informed decisions (especially in high-stake domains) Sachan et al. (2020); Lertvittayakumjorn et al. (2021), better understand and hence gain trust of the AI systems Pu and Chen (2006); Toreini et al. (2020), and even learn from the AI systems to accomplish tasks more successfully Mac Aodha et al. (2018); Lai et al. (2020). For AI system designers and developers, explanations allow them to identify the problems and weaknesses of the systems Krause et al. (2016); Han et al. (2020), calibrate the confidence of the systems Zhang et al. (2020b), and improve them accordingly Kulesza et al. (2015); Lertvittayakumjorn and Toni (2021).
@@
Explainability is particularly desirable for evaluation metrics. Sai et al. (2022) suggest that explainable NLG metrics should focus on providing more information than just a single score (such as fluency or adequacy). Celikyilmaz et al. (2020) stress the need for explainable evaluation metrics to spot system quality issues and to achieve a higher trust in the evaluation of NLG systems. Explanations indeed play a vital role in building trust for new evaluation metrics.222As an illustrating example, Moosavi et al. (2021) distrust using the metric BERTScore Zhang et al. (2020a) applied to a novel text generation task, as it assigns a score of 0.74 to a nonsensical output, which could be taken as an unreasonably high value. While BERTScore may indeed be unsuitable for their novel task, the score of 0.74 is meaningless here, as evaluation metrics may have arbitrary ranges. In fact, BERTScore typically has a particularly narrow range, so a score of 0.74 even for bad outputs may not be surprising (BERTScore provides a scaling method that solves these issues to some degree). Explainability techniques would be very helpful in preventing such misunderstandings, e.g., by showing that the features BERTScore attends to align with human judgement. For instance, if explanations of the scores align closely with human reasoning and faithfully reflect a metric’s internal decision process, this metric may be more likely to gain acceptance within the research community (see §2.2 for a definition of faithfulness and related terminology).
@@
By contrast, if faithfulness is given and the explanations are counter-intuitive, users and developers might lower their trust and be alerted to take additional actions, such as trying to improve the metrics using insights from the explanations or looking for alternative metrics that are more trustworthy. Furthermore, explainable metrics can be used for other purposes: for example, when a metric produces a low score for a given input, highlighted words (a widely used method for explanation, see §4) in the input are natural candidates for manual post-editing (when working with reference-free metrics, see §2.1).
@@
This concept paper aims at providing a systematic overview of the existing efforts in explainable MT evaluation metrics and an outlook for promising future research directions. The main focus of this work lies on explainable evaluation metrics for MT, although many of our observations and discussions can likely be adapted to other NLG tasks.
@@
§5 Future Work: An analysis of underexplored research directions, and a collection of exemplary methods from other NLP domains to illustrate potential future paths. Here, we also discuss the usage of recent large language models (LLMs), like ChatGPT OpenAI (2023a) and GPT4 OpenAI (2023b).
With this work, we aim to solidify the field of explainable MT metrics and provide guidance for researchers, e.g., metric developers, who aim to better understand and explain MT metrics, as well as users that want to employ explainability techniques to explain metric outputs. In the long term, we envision that our work will aid the development of improved MT metrics and thereby help to improve the quality of machine translations. Potentially, explanations can also aid other use cases such as translation selection and semi-automatic labeling (see §2.3). We also hope that our work will more generally inspire explainable NLG metric design.
@@ 2.1 Machine Translation Evaluation Metrics
MT metrics grade machine translated content, the hypothesis, based on ground truth data. Metrics can be categorized along several dimensions (e.g., Sai et al., 2022; Celikyilmaz et al., 2020). Here, we focus on a minimal specification that we will leverage later. A summary is shown in Table 1. Some parts of this section refer to explainability (see definitions in §2.2).
@@
Input type | Whether source, reference translation or both are used as ground truth for comparison |
@@
Granularity | At which level a metric operates: Word-level, sentence-level, document-level |
Quality aspect | What a metric measures: Adequacy, fluency, etc. |
Learning objective | How a metric is induced: Regression, ranking, etc. |
@@
Input type We call a metric reference-based if it requires one or multiple human reference translations to compare with the hypothesis as ground truth, which can be seen as a form of supervision (e.g., Papineni et al., 2002; Zhao et al., 2019; Rei et al., 2022a). Reference-free metrics do not require a reference translation to grade the hypothesis (e.g., Zhao et al., 2020; Ranasinghe et al., 2020a; Belouadi and Eger, 2023). Instead, they directly compare the source to the hypothesis. In the literature, reference-free MT evaluation is sometimes also referred to as “reference-less” (e.g., Mathur et al., 2020) or “quality estimation” (e.g., Zerva et al., 2022). This dimension is important for explainable MT evaluation, as explainability techniques might need to consider which resources are available.
@@
Granularity Translation quality can be evaluated at different levels of granularity: word-level, sentence-level and document-level. The majority of metrics for MT return a single sentence-level score for each input sentence (e.g., Zhao et al., 2019; Zhang et al., 2020a). Beyond individual sentences, a metric may also score whole documents (multiple sentences) (Jiang et al., 2022; Zhao et al., 2023). In the MT community, metrics that evaluate translations at the word-level (for example, whether individual words are correct or not) are also common (Turchi et al., 2014; Shenoy et al., 2021). Recently, there has been a tendency to compare metrics to human scores derived from fine-grained error analysis as they tend to correspond better to human professional translators (Freitag et al., 2021a). Metrics of higher granularity (e.g., word-level) provide more explanatory value than those of lower granularity, as they provide more details on which parts of the input might be translated incorrectly (e.g., Leiter, 2021; Fomicheva et al., 2021, 2022). Due to their prevalence, most techniques we describe in §4 target the explainability of metrics evaluated on sentence-level.
@@
Quality aspect This refers to the properties a metric evaluates and measures, for example, human assigned adequacy (e.g., via direct assessment, see Graham et al. (2016), obtained from crowd-workers), fluency, or other aspects (such as relevance, informativeness, or correctness, mostly in other NLG fields) (e.g., Sai et al., 2022; Celikyilmaz et al., 2020). We use this dimension in §4, as metrics that report multiple quality aspects provide more information on specific translations, i.e., they have more explanatory value for potential users.
@@
Learning objective Yuan et al. (2021) describe a further distinction based on the task a metric is designed to solve (or how a metric is implemented) to achieve high correlation with human judgments. They identify the three tasks below.333Chen and Eger (2023) argue for metrics induced from natural language inference (NLI) as a conceptual framework.
@@
Supervised Regression/Ranking: These models use supervised regression to train a metric to predict continuous values that directly correlate with human judgments from hypothesis and ground truth tokens. Alternatively, a metric may be trained to rank hypothesis sentences. For example, Rei et al. (2020) propose a ranking setup in which they minimize a loss that becomes smaller the greater the distance between embedding representations of the two sentences become. Especially in the MT community, metrics trained on human scores (e.g., COMET ;Rei et al. 2022a) have become more and more dominant, while Belouadi and Eger (2023) argue that fully unsupervised metrics, i.e., metrics that were constructed without reference sentences, parallel data and human scores, are more widely applicable.
@@
Text Generation: Some metrics use the probability that a sentence is created by a text generation system with a paraphrasing or translation task as sentence-level score. For example, the probability of translating a source sentence into a hypothesis sentence with an MT model can be used as score. Yuan et al. (2021) place the metric PRISM (Thompson and Post, 2020) and their own metric BARTScore in this category. In contrast, some recent metrics are based on in-context learning and directly generate scores (e.g. Kocmi and Federmann, 2023a; Lu et al., 2023b).
@@
Sentence-level MT metrics also have multiple use cases; see Figure 1 (in the following, numbers written in brackets correspond to this figure). (1) One use case is to rate a single translated text, for example, in a scenario where a human user wants to check their own human444Note, however, that humans might make different types of translation errors (Specia et al., 2018). Therefore, metric performance first needs to be evaluated for this use case. translation or a translation they received by an MT model before using it in a downstream task (e.g., Murgolo et al., 2022; Zouhar et al., 2023). This setup can also be used to filter parallel corpora for high-quality training data of new MT models (e.g., Ramos et al., 2023; Peter et al., 2023). (2) A second, common use case is to rate a large number of translations of different texts by one MT system to obtain a system-level score (for example by averaging sentence-level scores) (e.g., Freitag et al., 2022). Then, different MT systems can be compared their system-level scores that were computed on the same corpus. (3) Another use case is to employ metrics in the training process of MT models, e.g., using reinforcement learning (e.g., Wu et al., 2018) or optimizing by computing derivatives with Gumbel-softmax (e.g., Jauregi Unanue et al., 2021). (4) The fourth use case is to use the metrics in the decoding process of machine translation systems for reranking or with minimum Bayes risk decoding Fernandes et al. (2022a).
The use cases can have different audiences and explainability might play a different role to them (see 2.2). For example, anyone might be a user that wants to rate a translation, but usually only certain user groups, such as translation experts or MT experts, want to rate a translation model.
@@ 2.2 Explainability
Explainability & interpretability Prior work has noted that the terms interpretability and explainability vary in definition and are sometimes used ambiguously throughout the literature (Barredo Arrieta et al., 2020; Jacovi and Goldberg, 2020; Vilone and Longo, 2021; Freiesleben and König, 2023). In this paper, we follow the notion of passive interpretability and active explainability (for example, Barredo Arrieta et al., 2020). In particular, we adapt the definition of Barredo Arrieta et al. (2020) as follows:
@@
This definition varies from Barredo Arrieta et al. (2020) in two aspects: (1) It includes external explainers (also called post-hoc methods) and (2) it includes explanations of the metric’s output. We note that (2) is different from explaining the metric’s decision process, as output explanations may not provide insights on how the metric works. Still, we include (2) in our explainability definition for two reasons. One is that there have already been multiple shared tasks and other works on metric explainability that only evaluate how plausible explanations explain the output of a metric (e.g., Fomicheva et al., 2020, 2021; Zerva et al., 2022; Xu et al., 2023). We want to accommodate these works in our definition. The other reason is that there are use cases for such explanations of the output (specifically for the goals of making metrics more accessible and semi-automatic labeling; see §2.3).
@@
In contrast, a metric is interpretable if it can be passively understood by humans (for example, decision trees and k-nearest neighbors) (Barredo Arrieta et al., 2020). Modern MT metrics are mostly based on embeddings and black box language models, making them non-interpretable. The papers we categorize in this work generate explanations for MT metrics and their outputs, i.e., they consider explainability.
Faithfulness & plausibility An explanation is faithful if it accurately represents the reasoning process underlying a model’s prediction(s) (Jacovi and Goldberg, 2020). In contrast, an explanation is plausible if it convinces its audience that a model’s prediction is correct (Jacovi and Goldberg, 2021), i.e., if it aligns with human judgment. We illustrate the properties with the following example:
Source: Things became complicated
Hypothesis: Die Dinge bekamen kompliziert
Metric(Source, Hypothesis) = 0.8
Explanation 1: The word became is wrongly translated into bekamen instead of wurden. However, the meaning of the translation can easily be inferred. Therefore, on a scale of 0 to 1, a score of 0.8 is given.
Explanation 2: The words Die Dinge have 8 letters. Therefore, a score of 0.8 is assigned.
In this example, the English source sentence was (wrongly) translated into the German hypothesis sentence. A metric grades the translation with a score of 0.8 and two explanations (here, natural language explanations) are provided. Comparing these explanations, explanation 1 is a more plausible description (to translation experts) of why the score of 0.8 was assigned, as it bases the score on translation errors. However, which of the explanations describes the metric more faithfully cannot be inferred from the text. For example, internally the metric could be relying on word lengths, making explanation 2 more faithful, even though it is linguistically non-sensical.
When explainability techniques are used for metrics, faithfulness is always desirable and often, depending on the goal of the explanations, required. Plausibility (to translation experts), on the other hand, is required for some goals and not important for others (see §2.3).
As a side note, “complete faithfulness” of an explainability technique may never be established because it would require capturing the full decision process, e.g., of neural networks, in human understandable descriptions. For example, Jacovi and Goldberg (2020) deem it as “a unicorn which will likely never be found”.
Fairness & robustness We also categorize methods that test the robustness and fairness of a model (here, a metric) as explainability techniques. This is because these tests produce explanations that will help certain audiences to understand a model’s behavior better Barredo Arrieta et al. (2020).
Here, robustness measures how well a model can address execution errors, erroneous and malicious inputs and unseen data (e.g., Li et al., 2023a).
In contrast, Li et al. (2023a) define fairness as the goal that a model should not have biases that result in unfair treatment for certain groups of people.555There are many related and partly conflicting definitions and terms in explainable AI research (e.g., Vilone and Longo, 2021) and the given definitions are not yet firmly established in the literature. For example, Sharma et al. (2020) define explainability, robustness and fairness as exclusive properties. The main focus of our work is to build a taxonomy of methods for explainable MT metrics and the chosen definitions allow us to reasonably relate existing work.
In the MT metrics literature, this overlap in terminology is apparent for techniques that check how robust metrics are to input perturbations. Here, some works frame these as explainability techniques (e.g., Sai et al., 2021; Karpinska et al., 2022), while others frame them as robustness tests (e.g., He et al., 2023; Chen and Eger, 2023); see §4.
Robustness could be measured coarse-grainedly with a single score, such as the minimal number of adversarial perturbations required to successfully attack a metric, or fine-grainedly, by using predefined perturbation templates (e.g., Sai et al., 2021) or attackers that will attack with perturbations of specific properties (e.g., Chen and Eger, 2023). While a single score can also explain whether metric behavior is robust, susceptibility scores to various properties allow to assess metrics in greater detail. The same is true for fairness (Sun et al., 2022).
@@
understanding and model-understanding (e.g., Gehrmann et al., 2020; Holzinger et al., 2022).
@@
explaining general properties of a model that steer its decisions.
For example, a decision-understanding method could show that a metric assigned a score of 0.5, because word xyz was translated incorrectly. In contrast, a model-understanding method could show that a metric will always return a score of 1, given certain conditions.
@@
Often, decision-understanding is also referred to as local and model-understanding as global (e.g., Lipton, 2016; Guidotti et al., 2018; Doshi-Velez and Kim, 2017; Danilevsky et al., 2020).